OpenStudy (kitkat1):

Consider the following pair of equations: y = x + 4 y = –2x – 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (kitkat1):

I will give medal to the person that answers

OpenStudy (kitkat1):

Please Help me

OpenStudy (anonymous):

Okay. Lets see here. What do you mean by substitution?

OpenStudy (kitkat1):

U have to somehow substitute into the second equation

OpenStudy (anonymous):

okay....

OpenStudy (kitkat1):

I think like this y= -2(x+4) -2

OpenStudy (anonymous):

Okay. So... we need to find X. Lets make a balanced equation. y+2=-2x-8. y+10=-2x

OpenStudy (anonymous):

will that help?

OpenStudy (kitkat1):

I just dont get what to do next

OpenStudy (kitkat1):

like the steps to finding the answer

OpenStudy (kitkat1):

What Do i do next?

OpenStudy (kitkat1):

Someone Please Help will give medal!

OpenStudy (wolf1728):

Okay let's try this A) y = x + 4 B) y = –2x – 2 Let's put equation A into B x+4 = –2x – 2 3x = -6 x = ?

OpenStudy (kitkat1):

= -2?

OpenStudy (wolf1728):

Yes that's right.

OpenStudy (kitkat1):

Okay now I plug that in the first one right?

OpenStudy (kitkat1):

y = -2 + 4 ?

OpenStudy (kitkat1):

So the solution is (2, -2)

OpenStudy (kitkat1):

I havent seen how u got the 3x= -6 can u show me?

OpenStudy (wolf1728):

x = -2 and y = 2 so you would write the solutions as you would an (x,y) point or (-2, 2)

OpenStudy (wolf1728):

Okay I'll just recopy the answer I previously gave: A) y = x + 4 B) y = –2x – 2 Let's put equation A into B x+4 = –2x – 2 3x = -6 x = ?

OpenStudy (kitkat1):

Yes I just dont get how u got 3x= -6

OpenStudy (wolf1728):

You do see how equation A goes into B right?

OpenStudy (kitkat1):

yes

OpenStudy (kitkat1):

but then how did u turn it to 3x = -6

OpenStudy (wolf1728):

Okay here's equation A into equation B x+4 = –2x – 2

OpenStudy (kitkat1):

Okay and what u did next to get 3x = -6

OpenStudy (wolf1728):

Add 2x to both sides of the equation

OpenStudy (kitkat1):

Like this x + 4 = -2x -2 +2x +2x

OpenStudy (wolf1728):

No, like this x+4 = –2x – 2 now adding 2x to both sides x+4 +2x = –2x – 2 +2x Can you simplify that?

OpenStudy (kitkat1):

okay 4 + 2 = 6 and -2+ 2= 0

OpenStudy (wolf1728):

No, you have to add or subtract the x terms x+4 +2x = –2x – 2 +2x 3x + 4 = -2x + 2x -2 3x + 4 = -2 You follow that?

OpenStudy (kitkat1):

okay got that so now we have 3x + 4 = -2

OpenStudy (kitkat1):

3x + 4 = -2 -4 -4 = 3x = -6

OpenStudy (wolf1728):

3x + 4 = -2 Now we add -4 to both sides 3x + 4 -4 = -2 -4 Can you simplify that?

OpenStudy (kitkat1):

yes 3x = -6

OpenStudy (kitkat1):

Thank you very much!

OpenStudy (wolf1728):

u r welcome :-)